OpenStudy (anonymous):

Absolute Equations - Is my textbook wrong? 3|x+2| -1 = 8 I did the problem along wit the book and we both go: x = 1 or x = -5 Then we checked it 3|1+2| -1 ? 8 For which we BOTH got 8 and 3|-5 + 2| -1 ? 8 However here they got 8 and I got -10 Am I wrong and if so how??

OpenStudy (anonymous):

You want to show that your book is wrong?? :P

OpenStudy (anonymous):

For, this you must be aware of MODULUS.. |..| this has some meaning..

OpenStudy (anonymous):

\[|x| = |-x| = x\]

OpenStudy (anonymous):

Whether the value inside the modulus is positive or negative, the final result after removing Modulus brackets, will be Positive..

OpenStudy (anonymous):

For example: \(|3| = 3\) and also \(|-3| = 3\)

OpenStudy (anonymous):

This I knoooow but this was the equation before the brackets came off 3|-3| -1 how can that possibly come out to positive 8 or even -8 it didn't even come out to negative 8 (which I already know is equal to positive 8) @waterineyes

OpenStudy (anonymous):

I did not get you.. :(

OpenStudy (anonymous):

See : \[3 |-5 +2| -1 \implies 3|-3| - 1 \implies 3 \color{green}{(3)} -1 = ??\]

OpenStudy (anonymous):

I did this: 3|-5 +2| -1 3|-3| -1 -9 -1 -10 So wait @waterineyes once I get to 3|-3|-1 the 3 becomes positive. Oooh wait I get it because the absolute value of -3 is 3 thank you!

OpenStudy (anonymous):

yep.. :)

OpenStudy (anonymous):

Use absolute value first and then multiply it with \(3\).. :)

OpenStudy (anonymous):

\(|x|\) = \(\begin{cases} x, && x > 0 \\-x, && x < 0\end{cases}\)

OpenStudy (anonymous):

So, whenever you deal with Mod, then you have to divide it into two parts, one greater than 0 and other less than 0.

OpenStudy (anonymous):

\(|x+2|\) = \(\begin{cases} (x+2), && (x+2) > 0 \\-(x+2) , && (x+2) < 0\end{cases}\)

OpenStudy (anonymous):

So, one time use (x+2) and solve the equation, you will get one value.. And next time : use -(x+2) and solve, you will get another value. :)

OpenStudy (anonymous):

That's the best explanation thank you. :)

OpenStudy (anonymous):

you are welcome dear.. :)